+8 (b) s(+16) 2s3 3s +6s+4 (s2 +4)(s2 +2s+2) (c) +3s + (d) 2 2 1 s-2s +s (e) (s+2)e (f 2s +1 22-6s3 (g) -3s +2 2(s22r4) (24) find. inverse Laplace transfer show step buy step soultion
Answer #1

S2+8 13 S+C A s(s+16) AS416A gs+ cs On Conn payma coeffiuenfsa s s, Cemslant de t Ay o 14A- A+8-1B -A s2+9 B.S+C S(s2+ On app-AS A+R+2 C=2 A,R VAlue ユーc+2-2り +2C 3 c -2D 1 2) DEI C AI 8=0 23 + s244 2s+2) る+3てナS S+ st4 $2+2) S+ verse laphce cosal Buy1-3(S+3a fnphce in Verse &C -3e Casht snhat + eu coe hat btcashat (4b s2-s+ s2-s+ s322 (G-S ACs1-25+B Cs +cs &(s ST - con7aySt2 S+ (St1) (S41)2 (s+1) St On apply ing verse phce - ettet S+2 (-+2 +4-1)t- Now act-6) bs then Sta 2s-6545 a darerfenchin C2 Cs +4) ASH CS+D (24 AS+B ) +CS+D (sty 4 A+C = f E - = 0 2 Cs22s+4 ) 4S 12 Verre lapce Buy Rydd sing t 4t cos t 654)2 (+56 B(hifhitheren 4 - 25 BS+C first tete AS16A + BS cs S2+8 conParing A t&= C-O 6A- SLT16 l여hce mYexse On Icos4t By gecnd shiftin